If 7182 roses were killed in last night's frost and this was 2 5/8 as many as the night before how many roses were killed the night before?
Answer by stanbon(75887)

If 7182 roses were killed in last night's frost and this was 2 5/8 as many as the night before how many roses were killed the night before?
---
Let the number killed the night before be "x":
Equation:
(2 5/8)x = 7182
---
(21/8)x = 7182
---
x = (8/21)7182
----
x = 2736
===============
Cheers,
Stan H.
==================